# Position Coupling: Leveraging Task Structure for Improved Length Generalization of Transformers

### v1

@@ F.3.1 Token Embedding
Dimension 1 (num).   For a number token ($0,\ldots,9$), we put itself into the dimension num. For the other tokens (×,=,$\times$,=,\\$× , = , \$), we put 00.
Dimension 2 (full_ones).   We put $1$ everywhere in this dimension.
Dimension 3 (is_bos).   For a token ‘\$currency-dollar\\$\$’, we put $1$ into the dimension is_bos. Otherwise, we put 00.
Dimension 4 (is_mul).   For a token ‘$\times$’, we put $1$ into the dimension is_mul. Otherwise, we put 00.
Dimension 5 (is_equal).   For a token ‘$=$’, we put $1$ into the dimension is_equal. Otherwise, we put 00.
F.3.2 Coupled Position IDs and Position Embedding
@@
F.4 Construction Idea
Here, we provide an example that demonstrates how we construct the $N\times 2$ multiplication. Consider the calculation $7595\times 79=600005$. While a typical method for computing such a multiplication is illustrated in Table 29, we consider an alternative approach, as shown in Table 30. In this method, we pair the digits from the first and second operands at each step where the sum of their digit positions is the same, and then calculate the sum of the pairwise products. For example, the number $116$ in Table 30 is generated by ${\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}9}\times{\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}5}\times{\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}7}$, and the number $108$ is generated by ${\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}5}\times{\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}\definecolor[named]{pgfstrokecolor}{rgb}{.094,.455,.804}9}\times{\color[rgb]{.804,.15,.15}\definecolor[named]{pgfstrokecolor}{rgb}{.804,.15,.15}7}$, where blue indicates numbers from the first operand and red indicates numbers from the second operand. The main reason for considering such a method is to provide a clearer intuition for determining which numbers from each operand we should attend to when predicting the next token.
@@ F.4 Construction Idea
[…]
@@
F.5 Transformer Block 1 — Causal Attention Layer
To implement the concept introduced in Section F.4, it is essential to design a Transformer block capable of generating an embedding matrix depicted in the left-most table of Table 31. The goal of the first Transformer block is to fill is_op2_one ($6$-th dimension) and is_op2_ten ($7$-th dimension) by $1$ if the token corresponds to the ones or tens digit of the second operand, respectively, and 0 otherwise. These two dimensions enable the filling of op2_one ($8$-th dimension) and op2_ten ($9$-th dimension) at the second Transformer block. Furthermore, we will fill mask ($34$-th dimension) in the first block, which will serve as a base for filling op1_shift0 to op1_shift4 in the second block. Thus, we currently have 3 objectives, each of which will be addressed by an individual head.
F.5.1 Attention Head 1: Detecting the Ones Digit of the Second Operand
@@
By these, the attention score matrix ${\bm{C}}^{(1)}_{1}:=({\bm{K}}^{(1)}_{1}{\bm{X}}^{(0)})^{\top}{\bm{Q}}^{(1)}_{1}{\bm{X}}^{(0)}$ and the attention matrix ${\bm{A}}^{(1)}_{1}:={\tt softmax}({\bm{C}}^{(1)}_{1})\in\mathbb{R}^{N\times N}$ can be obtained. We provide the example of ${\bm{A}}_{1}^{(1)}$ in Table 34.
@@
F.5.2 Attention Head 2: Detecting the Tens Digit of the Second Operand
In the previous head, we set the dimension is_op2_one ($6$-th dimension) to a one-hot row vector, where $1$ is placed only in the token corresponding to the ones digit of the second operand. The objective of Attention head 2 is to fill the dimension is_op2_ten ($7$-th dimension) in a similar manner to is_op2_one, but with $1$ placed only in the tens digit of the second operand.
@@ F.5.2 Attention Head 2: Detecting the Tens Digit of the Second Operand
F.5.3 Attention Head 3: Position Masking
@@ F.5.4 Residual Connection
F.6 Transformer Block 1 — Token-wise Feed-forward Layer
@@
Construction for pos_2_mask.   Given a vector ${\bm{y}}=[{\bm{y}}_{j}]_{j=1}^{d}\in\mathbb{R}^{d}$, define functions $g_{l},h_{l}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ for every $j\in[P]$ as
@@
$h_{l}({\bm{y}}):=-{\bm{y}}_{\textsc{pos\_2},l}-2{\bm{y}}_{\textsc{mask}},$ (96)
@@
Table 48: Example output of FFN layer at the first Transformer block, continuing from Table 47. Omitted entries are filled with 00.
@@ F.6.1 Residual Connection
F.7 Transformer Block 2 — Causal Attention Layer
@@
F.7.1 Attention Head 1: Copying the Ones Digit of the Second Operand
@@
F.7.2 Attention Head 2: Copying the Tens Digit of the Second Operand
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ F.7.8 Residual Connection
F.8 Transformer Block 2 — Token-wise Feed-forward Layer
As mentioned in the theorem statement below, we allow ${\tt FF}_{2}$ to be a multi-layer MLP. Our ultimate goal is to fill the dimensions prod and is_eos with appropriate values. The dimensions result1 to result4, pre_prod, and pre_carry serve as temporary memories for storing intermediate values, which will help us achieve our ultimate goal. Our construction involves sequentially stacking the MLP networks step-by-step to generate each of these temporary values.
@@ F.8 Transformer Block 2 — Token-wise Feed-forward Layer
Recalling Section F.4, we construct the first MLP network by utilizing eight instances of the function $f$ in parallel as follows:
@@
Recalling Section F.4, we construct the second MLP network on top of the first MLP network, by utilizing 2 instances of each of the function $g_{1}$, $g_{2}$, and $g_{3}$ in parallel as follows:
@@
Similarly to the previous step, each network can be implemented with 280, 20, and 10 hidden nodes. Recalling Section F.4, we construct the third MLP network, on top of the second MLP network, by
$\textsc{prod}=\begin{pmatrix}q_{0}(h^{\prime}(\textsc{pre\_prod}+h(\textsc{pre\_carry},\textsc{num})))\\ q_{1}(h^{\prime}(\textsc{pre\_prod}+h(\textsc{pre\_carry},\textsc{num})))\\ \vdots\\ q_{9}(h^{\prime}(\textsc{pre\_prod}+h(\textsc{pre\_carry},\textsc{num})))\\ \end{pmatrix}\in\mathbb{R}^{10}$
@@ F.8.1 Residual Connection
F.9 Decoding Function
As mentioned in Appendix C, the decoding function performs a linear readout (with a weight matrix ${\bm{W}}_{\rm out}\in\mathbb{R}^{\left|{\mathcal{V}}\right|\times d}$) and a (token-wise) arg-max operation. That is,

### v2

@@ G.3.1 Token Embedding
Dimension 1 (num).
For a number token ($0,\ldots,9$), we put itself into the dimension num. For the other tokens (×,=,$\times$,=,\\$× , = , \$), we put 00.
Dimension 2 (full_ones).
We put $1$ everywhere in this dimension.
Dimension 3 (is_bos).
For a special token ‘\$currency-dollar\\$\$’, we put $1$ into the dimension is_bos. Otherwise, we put 00.
Dimension 4 (is_mul).
For a special token ‘$\times$’, we put $1$ into the dimension is_mul. Otherwise, we put 00.
Dimension 5 (is_equal).
For a special token ‘$=$’, we put $1$ into the dimension is_equal. Otherwise, we put 00.
G.3.2 Coupled Position IDs and Position Embedding
@@
G.4 Construction Idea
Here, we provide an example that demonstrates how we construct the $N\times 2$ multiplication. Consider the calculation $7595\times 79=600005$. While a typical method for computing such a multiplication is illustrated in Table 29, we consider an alternative approach, as shown in Table 30. In this method, we pair the digits from the first and second operands at each step where the sum of their digit positions is the same, and then calculate the sum of the pairwise products. For example, the number $116$ in Table 30 is generated by ${\color[rgb]{.094,.455,.804}9}\times{\color[rgb]{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}5}\times{\color[rgb]{.804,.15,.15}7}$, and the number $108$ is generated by ${\color[rgb]{.094,.455,.804}5}\times{\color[rgb]{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}9}\times{\color[rgb]{.804,.15,.15}7}$, where blue indicates numbers from the first operand and red indicates numbers from the second operand. The main reason for considering such a method is to provide a clearer intuition for determining which numbers from each operand we should attend to when predicting the next token.
@@ G.4 Construction Idea
  • $49={\color[rgb]{.094,.455,.804}0}\times{\color[rgb]{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}7}\times{\color[rgb]{.804,.15,.15}7}$,
  • $98={\color[rgb]{.094,.455,.804}7}\times{\color[rgb]{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}5}+{\color[rgb]{.804,.15,.15}7}$,
  • $108={\color[rgb]{.094,.455,.804}5}\times{\color[rgb]{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}9}+{\color[rgb]{.804,.15,.15}7}$,
  • $116={\color[rgb]{.094,.455,.804}9}\times{\color[rgb]{.804,.15,.15}9}+{\color[rgb]{.094,.455,.804}5}+{\color[rgb]{.804,.15,.15}7}$.
@@
G.5 Transformer Block 1 — Causal Attention Layer
To implement the concept introduced in Section G.4, it is essential to design a Transformer block capable of generating an embedding matrix depicted in the left-most table of Table 31. The goal of the first Transformer block is to fill is_op2_one ($6$-th dimension) and is_op2_ten ($7$-th dimension) by $1$ if the token corresponds to the ones or tens digit of the second operand, respectively, and 0 otherwise. These two dimensions enable the filling of op2_one ($8$-th dimension) and op2_ten ($9$-th dimension) at the second Transformer block. Furthermore, we will fill mask ($34$-th dimension) in the first block, which will serve as a base for filling op1_shift0 to op1_shift4 in the second block. Thus, we currently have 3 objectives(is_op2_one, is_op2_ten, mask), each of which will be addressed by an individual head.
G.5.1 Attention Head 1: Detecting the Ones Digit of the Second Operand
@@
By these, the attention score matrix ${\bm{C}}^{(1)}_{1}:=({\bm{K}}^{(1)}_{1}{\bm{X}}^{(0)})^{\top}{\bm{Q}}^{(1)}_{1}{\bm{X}}^{(0)}$ and the attention matrix ${\bm{A}}^{(1)}_{1}:={\tt softmax}({\bm{C}}^{(1)}_{1})\in\mathbb{R}^{N\times N}$ can be obtained. We provide the example of ${\bm{A}}_{1}^{(1)}$ in Table 34. Specifically, an entry in ${\bm{A}}_{1}^{(1)}$ is non-zero if and only if the inner product between the query and key vectors equals $MP$.
@@
G.5.2 Attention Head 2: Detecting the Tens Digit of the Second Operand
In the previous head, we set the dimension is_op2_one ($6$-th dimension) to a one-hot row vector, where $1$ is placed only in the token corresponding to the ones digit of the second operand. The objective of Attention head 2 is to fill the dimension is_op2_ten ($7$-th dimension) similarly to is_op2_one, but with $1$ placed only in the tens digit of the second operand.
@@ G.5.2 Attention Head 2: Detecting the Tens Digit of the Second Operand
G.5.3 Attention Head 3: Position Masking
@@ G.5.4 Residual Connection
G.6 Transformer Block 1 — Token-wise Feed-forward Layer
@@
Construction for pos_2_mask.
Given a vector ${\bm{y}}=[{\bm{y}}_{j}]_{j=1}^{d}\in\mathbb{R}^{d}$, define functions $g_{l},h_{l}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ for every $j\in[P]$ as
@@
$h_{l}({\bm{y}}):=-{\bm{y}}_{\textsc{pos\_2},l}-2{\bm{y}}_{\textsc{mask}}$ (95)
@@
Table 48: Example output of FFN layer at the first Transformer block, continuing from Table 47.
@@ G.6.1 Residual Connection
G.7 Transformer Block 2 — Causal Attention Layer
@@
G.7.1 Attention Head 1: Copying the Ones Digit of the Second Operand
@@
G.7.2 Attention Head 2: Copying the Tens Digit of the Second Operand
@@
G.7.3 Attention Head 3: Copying the Appropriate Digit from the First Operand I
@@
G.7.4 Attention Head 4: Copying the Appropriate Digit from the First Operand II
@@
G.7.5 Attention Head 5: Copying the Appropriate Digit from the First Operand III
@@
G.7.6 Attention Head 6: Copying the Appropriate Digit from the First Operand IV
@@
G.7.7 Attention Head 7: Copying the Appropriate Digit from the First Operand V
@@ G.7.8 Residual Connection
G.8 Transformer Block 2 — Token-wise Feed-forward Layer
Our ultimate goal is to fill the dimensions prod and is_eos with appropriate values. The dimensions result1 to result4, pre_prod, and pre_carry serve as temporary memories for storing intermediate values, which will help us achieve our ultimate goal. Our construction involves sequentially stacking the MLP networks step-by-step to generate each of these temporary values. (As mentioned in the theorem statement below, we allow ${\tt FF}_{2}$ to be a multi-layer MLP.)
@@ G.8 Transformer Block 2 — Token-wise Feed-forward Layer
Recalling Section G.4, we construct the first MLP network by utilizing eight instances of the function $f$ in parallel as follows:
@@
Recalling Section G.4, we construct the second MLP network on top of the first MLP network, by utilizing 2 instances of each of the function $g_{1}$, $g_{2}$, and $g_{3}$ in parallel as follows:
@@
Similarly to the previous step, each network can be implemented with 280, 20, and 10 hidden nodes. Recalling Section G.4, we construct the third MLP network, on top of the second MLP network, by
  • $\textsc{prod}=\begin{pmatrix}q_{0}(h^{\prime}(\textsc{pre\_prod}+h(\textsc{pre\_carry},\textsc{num})))\\ q_{1}(h^{\prime}(\textsc{pre\_prod}+h(\textsc{pre\_carry},\textsc{num})))\\ \vdots\\ q_{9}(h^{\prime}(\textsc{pre\_prod}+h(\textsc{pre\_carry},\textsc{num})))\\ \end{pmatrix}\in\mathbb{R}^{10}$.
@@ G.8.1 Residual Connection
G.9 Decoding Function
As mentioned in Appendix D, the decoding function performs a linear readout (with a weight matrix ${\bm{W}}_{\rm out}\in\mathbb{R}^{\left|{\mathcal{V}}\right|\times d}$) and a (token-wise) arg-max operation. That is,